# The sphere complex of a locally finite graph

### v1

@@ 7.2 Nonseparating spheres and finite rank graphs
To define the notion of projection we will use, we need the following lemma which will imply the independence of the choice of representative of the sphere we are projecting.
@@
We can assume that $S$, and therefore $S^{\prime}$, are not disjoint from $K$ or equal to a component of $\partial K$. Thus such a $T$ as described in the statement of the lemma actually exists. Let us clarify that if the discs in $\partial K$ used to define $T$ are nested, to obtain $T$ as an embedded sphere we need to move some of the discs slightly into $K$. We can assume that this happens on some fixed small neighborhood of $\partial K$.
@@
The next lemma is stated in full generality, though we will only need it when $\Gamma$ is a finite rank graph.
@@
Our goal to replace $\gamma$ with a curve $\gamma^{\prime}$ which lies in $K$ that intersects $R$ in the same point, and nowhere else. Once this is done, the result follows. Let $U$ denote a complementary component of $K$, and suppose $a$ is a component of $\gamma\cap U$. If $S$ does not intersect $U$, it is clear we may modify $\gamma$ by replacing $a$ with another arc between the intersection points of $a$ with $\partial K$ with a small arc that lies entirely in $\partial K$. This new curve does not intersect $S$ at any new points.
@@
But now, by Lemma 7.9, the endpoints of $a$ are in the same component of $\partial U\setminus S$, so we replace $a$ with another arc as described above. Repeating this for all arcs of $\gamma$ that leave $K$, we obtain the curve $\gamma^{\prime}$ as desired. ∎
@@
where the first inequality is Lemma 2.16(b), the second inequality is Lemma 2.16(a), and the finally the third follows from the bi-Lipschitz lower bounds for $i$. As $\psi$ has positive translation length on $\mathcal{AF}_{n}$, it follows that the same is true on $\mathcal{S}_{ns}(M_{\Gamma})$.
Now suppose $\text{rk}(\Gamma)=1$. In this case, we take $K$ to be a full submanifold with $2$ boundary components. As discussed in Subsection 2.3, $\mathcal{S}_{ns}(K)$ is isomorphic to the real line with vertices at the integers. The action of $\text{Map}(M_{1,2})$ has elements with positive translation length. For example, one can consider maps which push one of the boundary spheres around a nontrivial loop. Thus it follows as above that there are elements of $\text{Map}(\Gamma)$ which have positive translation length for the action on $\mathcal{S}_{ns}(M_{\Gamma})$. ∎
@@
Recall that the corresponding conjecture for the full sphere complex $\mathcal{S}(M_{n,s})$ is known to be false if $s\geq 3$, as in this case $\mathcal{S}(M_{n,s})$ has finite diameter, and false when $n\geq 4$ and $s=2$ as $\mathcal{S}(M_{n,s})$ is not hyperbolic. See the discussion in Subsection 2.3.
@@ 7.3 Outer space for locally finite graphs
For point stabilizers, note that any $\mathcal{S}\in\mathbb{O}(\Gamma)$ divides $M_{\Gamma}$ into finitely many components, all of which are simply connected and with finitely many sphere boundary components. For the compact components, the only nontrivial elements are those which exchange boundary spheres, and there are only finitely many ways to do this, with some restrictions coming from the weight of the spheres. For the noncompact components, the same is true, except one can also take homeomorphisms of the end space. One can assume these homeomorphisms are all supported outside of a neighborhood of the boundary components of the component. The result follows, as the stabilizer of $\mathcal{S}$ will consist of a semidirect products of subgroups of the groups just described. ∎
@@ 8.2 Translatable graphs
Are there any examples of graphs $\Gamma$ so that $\mathcal{TD}(B)$ is quasi-isometrically embeds into to $\mathcal{TC}(\Sigma)$ or $\mathcal{TS}(M_{\Gamma})$, and so that all three graphs have infinite diameter? What about examples where the inclusion maps are quasi-isometries?

### v2

@@ 7.2 Nonseparating spheres and finite rank graphs
To define the notion of projection we will use, we need the following lemma which will imply the independence of the choice of representative of the sphere we are projecting. Recall the notions of normal form and equivalence of sphere systems defined in Subsection 2.2.
@@
We can assume that $S$, and therefore $S^{\prime}$ nontrivially intersect $\partial K$, and thus such a $T$ as described in the statement of the lemma actually exists. Let us clarify that if the discs in $\partial K$ used to define $T$ are nested, to obtain $T$ as an embedded sphere we need to move some of the discs slightly into $K$. We can assume that this happens on some fixed small neighborhood of $\partial K$.
@@
The next lemma is stated in full generality, though we will only need it when $\Gamma$ is a finite rank graph. We note the following essential fact that is used in the following lemma. A sphere $S\subset K$ for $K$ a submanifold of $M_{\Gamma}$ is nonseparating in $K$ if and only if there is a closed curve that lies in $K$ which intersects $S$ exactly one time.
@@
Our goal to replace $\gamma$ with a curve $\gamma^{\prime}$ which lies in $K$ that intersects $R$ in the same point, and nowhere else. Once this is done, the result follows, as any sphere which results from surgery using the component $R$ will be nonseparating by the fact mentioned before the lemma. Let $U$ denote a complementary component of $K$, and suppose $a$ is a component of $\gamma\cap U$. If $S$ does not intersect $U$, it is clear we may modify $\gamma$ by replacing $a$ with another arc between the intersection points of $a$ with $\partial K$ with a small arc that lies entirely in $\partial K$. This new curve does not intersect $S$ at any new points.
@@
But now, by Lemma 7.9, the endpoints of $a$ are in the same component of $\partial U\setminus S$, so we can replace $a$ with another arc as described above. Repeating this for all arcs of $\gamma$ that leave $K$, we obtain the curve $\gamma^{\prime}$ as desired. ∎
@@
where the first inequality is Lemma 2.16(b), the second inequality is Lemma 2.16(a), and the third follows from the bi-Lipschitz lower bounds for $i$ established in the first set of inequalities above. As $\psi$ has positive translation length on $\mathcal{AF}_{n}$, it follows that the same is true on $\mathcal{S}_{ns}(M_{\Gamma})$.
Now suppose $\text{rk}(\Gamma)=1$. In this case, we take $K$ to be a full submanifold with $2$ boundary components. As discussed in Subsection 2.3, $\mathcal{S}_{ns}(K)$ is isomorphic to the real line with vertices at the integers. The action of $\text{Map}(M_{1,2})$ has elements with positive translation length. For example, one can consider maps which push one of the boundary spheres around a nontrivial loop. Consider such a map $\psi$. Then, as
$d_{\mathcal{S}_{ns}(K)}(\psi^{m}(S),S)\\ \leq 3d_{\mathcal{S}_{ns}(M_{\Gamma})}(\psi^{m}(S),S)$
for any $S\in\mathcal{S}(K)$, it follows that there are elements of $\text{Map}(\Gamma)$ which have positive translation length for the action on $\mathcal{S}_{ns}(M_{\Gamma})$. ∎
@@
Recall that the corresponding conjecture for the full sphere complex $\mathcal{S}(M_{n,s})$ is trivial if $s\geq 3$, as in this case $\mathcal{S}(M_{n,s})$ has finite diameter, and false when $n\geq 4$ and $s=2$ as $\mathcal{S}(M_{n,s})$ is not hyperbolic. See the discussion in Subsection 2.3.
@@ 7.3 Outer space for locally finite graphs
Let $\Gamma$ be a locally finite graph with finite positive rank. Then $\mathbb{O}(\Gamma)$ is contractible. The point stabilizers of the action of $\mathrm{PMap}(\Gamma)$ on $\mathbb{O}(\Gamma)$ are subgroups are finite.
@@
For point stabilizers of elements pf $\mathbb{O}(\Gamma)$ in $\text{PMap}(\Gamma)$, note that any $\mathcal{S}\in\mathbb{O}(\Gamma)$ divides $M_{\Gamma}$ into finitely many components, all of which are simply connected and with finitely many sphere boundary components. By Theorems 1.1 and 2.1, the subgroup of the stabilizer of $\mathcal{S}$ which stabilizes each complementary component is finite, as by the theorems, diffeomorphisms on these components are determined up to isotopy and sphere twists by how they act on the boundary spheres and ends of the component. But by assumption, as we are working in the pure mapping class group, the ends are fixed, so even noncompact complementary components of $\mathcal{S}$ contribute only finitely many elements. As there are only finitely many complementary components, it follows that the stabilizer of $\mathcal{S}$ is also finite. ∎
One can also consider stabilizers of the action of $\text{Map}(\Gamma)$ on $\mathbb{O}(\Gamma)$. Typically, these will be infinite as they will include mapping classes which stabilize some noncompact component of the complement of the sphere system, but acts by a nontrivial homeomorphism on the space of ends of the component.
@@ 8.2 Translatable graphs
Are there any examples of translatable graphs $\Gamma$ so that $\mathcal{TD}(B)$ is quasi-isometrically embeds into to $\mathcal{TC}(\Sigma)$ or $\mathcal{TS}(M_{\Gamma})$, and so that all three graphs have infinite diameter? What about examples where the inclusion maps are quasi-isometries?